OpenStudy (yoyo):

1)n-3+4>7 2)3-6n-4<17(negative 4 i think?) 3)06 6) 8(1-4x)>40 7)5x-7(x+1)>-9 PLEASE SHOW ALL WORK THANK YOU.

OpenStudy (anonymous):

n+1>7 n>6 -6n<18 n>-3 when dividing, you have to flip the less then or greater then sine in the opposite direction. 0 < n + 5 -5<n -5x<5 x > -1 flip sine 2m >6 m > 3 8 -32x > 40 32x > 32 x > 1 5x-7x+7>-9 -2x+7>-9 -2x>-16 x < 8

OpenStudy (mercy):

n-3+4>7=n>6 3-6n-4<17=-6n-1<17=-6n<18=3n>-6 0<n-1+6=0<n+5=-5<n -3x-2x<5=-5x<5=x>-1 2+2m-2>6=m>3 1-4x>40=-4x>39=x<39/4 5x-(7x+1)>-9=-2x+1>-9=-2x>-8=x<4